# Noema formIng Cluster survEy (NICE): A Census of Star Formation and Cold Gas Properties in Massive protoclusters at 1.5 ¡ ¡ 4

### v2

@@
Massive protoclusters at , the peak of the cosmic star formation history, are key to understanding the formation mechanisms of massive galaxies in today’s clusters. However, studies of protoclusters at these high redshifts remain limited, primarily due to small sample sizes and heterogeneous selection criteria. In this work, we conduct a systematic investigation of the star formation and cold gas properties of member galaxies of eight massive protoclusters in the COSMOS field, using the statistical and homogeneously selected sample from the Noema formIng Cluster survEy (NICE). Our analysis reveals a steep increase in the star formation rates per halo mass () with redshifts in these intensively star-forming protoclusters, reaching values one to two orders of magnitude higher than those observed in the field at ¿ 2. We further show that, instead of an enhancement of starbursts, this increase is largely driven by the concentration of massive and gas-rich star-forming galaxies in the protocluster cores. The member galaxies still generally follow the same star formation main sequence as in the field, with a moderate enhancement at the low mass end. Notably, the most massive protocluster galaxies ( ¿ 81010M⊙) exhibit higher and than their field counterparts, while remaining on the star forming main sequence. These gas-rich, massive, and star-forming galaxies are predominantly concentrated in the protocluster cores and are likely progenitors of massive ellipticals in the center of today’s clusters. These results suggest that the formation of massive galaxies in such environments is sustained by substantial gas reservoirs, which support persistent star formation and drive early mass assembly in forming cluster cores.
@@
The environmental influence on the progenitors of massive galaxies in local clusters is critical for understanding the formation of these galaxies (Peng et al., 2010; Behroozi et al., 2019; van der Burg et al., 2020) . Over the past decade, extensive studies of massive protoclusters at the peak of the cosmic star formation history (), selected through various methods, have been inconclusive in producing an unambiguous, consistent picture of the environmental effects on star formation and cold gas properties of galaxies in these structures (e.g., Rudnick et al., 2017; Gómez-Guijarro et al., 2019; Jin et al., 2023, 2024; Bakx et al., 2024; Pensabene et al., 2024). Specifically, some structures show a prevalence of submillimeter galaxies (SMGs) or ultra-luminous infrared galaxies (LIRGs), suggesting enhanced star formation in their galaxies as compared to average field environments (e.g., Casey, 2016; Wang et al., 2016; Oteo et al., 2018). Other studies, primarily based on H emitters in protoclusters, report little to no differences in the star formation main sequence (SFMS) compared to their field counterparts (e.g, Valentino et al., 2015; Shimakawa et al., 2018; Pérez-Martínez et al., 2023). Molecular gas is the fuel for star formation. The well established star formation law (Kennicutt, 1998; Gao & Solomon, 2004) highlights the role of internal secular processes in the conversion of cold gas into stars. Additionally, a bimodal scenario has been proposed to describe the observed offset between starburst galaxies and normal star-forming galaxies, with starburst galaxies showing higher star formation efficiency (SFE), which may result from a higher merger rate, or a more compact distribution of molecular gas, etc., (Daddi et al., 2010; Sargent et al., 2014; Liu et al., 2019).
@@
In dense environments, increased interactions can compress gas, enhancing SFEs, while cold gas inflows through filaments connected to protoclusters can replenish gas reservoirs, sustaining high gas fractions. Deep CO observations of cluster galaxies 1.6 reveal a consistent or enhanced gas fraction compared to the field counterparts (Rudnick et al., 2017; Noble et al., 2017). In the Spiderweb protocluster at =2.15, Dannerbauer et al. (2017) identified a massive and extended CO(1-0) disk around the central radio galaxies. Follow-up analysis by Pérez-Martínez et al. (2025) revealed a decreasing gas fraction with increasing stellar mass, showing a particularly steep decline above log(/M⊙) = 10.5. In CL J1001, a starbusting protocluster at = 2.51, Wang et al. (2018) observed decreasing gas fractions towards the cluster center, suggesting the imminent formation of a passive core by 2. Tadaki et al. (2019) found a declining gas fraction and gas depletion time with increasing stellar mass in galaxies in three protoclusters traced by radio galaxies at , and suggested that gas accretion is accelerated in less massive galaxies while suppressed in more massive ones. At even higher redshifts ( 4), extreme starbusting protocluster cores such as SPT2349-56 and Distant Red Core (DRC) exhibit lower gas fractions compared to field SMGs (Long et al., 2020; Hill et al., 2022). Nevertheless, there remains a paucity of protoclusters at high redshift with deep, multi-wavelength coverage to comprehensively link the stellar and gas properties of member galaxies. Such observations are essential for revealing the physical mechanisms driving the formation of massive galaxies in these dense environments.
@@
The Noema formIng Cluster survEy (NICE, Zhou et al. 2024) is a 159 hours NOEMA Large Program (ID: M21AA, PIs: E. Daddi and T. Wang) targeting 48 massive protocluster candidates across multiple fields, complemented by a 40-hour ALMA program (ID: 2021.1.00815.S) focusing on 25 southern candidates. The primary goal of these programs is to spectroscopically confirm protoclusters at cosmic noon through CO emission from their member galaxies, significantly expanding the sample size in this field of research. Notably, eight protoclusters at 1.5 ¡ ¡ 4 have been confirmed in the COSMOS field (Sillassen et al., 2024). The COSMOS field provides extensive ancillary data across optical, near-infrared (NIR), far-infrared (FIR), submillimeter, and radio wavelengths (Weaver et al., 2022; Jin et al., 2018; Smolčić et al., 2017) which combined with the CO observations from the NICE survey, offer a unique opportunity to investigate the environmental impact on galaxy evolution.
@@
In this paper, we investigate the star formation activity and cold gas content of eight massive protoclusters in the COSMOS field, identified through the NICE survey (hereafter NICE-COSMOS). Our analysis focuses on the protoclusters’ central region (within one virial radius), as detailed in the following sections. We adopt a standard CDM cosmology with = 70 km s-1 Mpc-1, = 0.3, and = 0.7 and a Chabrier (2003) initial mass function.
@@ 2.1 NICE-COSMOS protoclusters
The eight protoclusters in NICE-COSMOS, along with their member identification, are detailed in Sillassen et al. (2024). Briefly, the members are selected with — - — ¡ 0.1(1 + ) within virial projected radius based on the COSMOS2020 catalog (Weaver et al., 2022). The NICE survey selects intensively star-forming massive protoclusters at 2 4, which are chosen as overdensities of high- massive galaxies traced by red IRAC sources in association with SPIRE-350m peakers that show intense star formation. We list the selection criteria here:
@@
, and are the Herschel/SPIRE fluxes at 250m, 350m, and 500m.
Fig. 8 shows the spatial distribution of all members including CO and continuum detections in the eight protoclusters in NICE-COSMOS, which are primarily concentrated at the cluster centers. For the following analysis, we focus on galaxies within one virial radius, 1 as defined in (Sillassen et al., 2024), as this area is also well covered by the NOEMA and ALMA observations. The virial radii range from 141 pkpc to 429 pkpc, and the number of member galaxies per structure varies between 13 and 31.
@@ 2.2 Stellar masses and star formation rates
We use stellar masses and star formation rates (SFRs) derived with LePhare (Arnouts et al., 2002) in COSMOS2020 (Weaver et al., 2022). For galaxies detected in the FIR, specifically those in the super-deblended catalog (Jin et al., 2018), we also explored various SFR tracers to obtain accurate measurements.
We first cross matched the member galaxies with the super-deblended catalog (Jin et al., 2018)with a separation smaller than 1. The matches have a median separation of 0.15. Notably, the 3GHz and 24m fluxes are measured directly at the galaxy positions, hence the risk of mis-identification or flux misattribution is substantially reduced. For galaxies detected at 3 GHz (Smolčić et al., 2017) with a signal-to-noise ratio (SNR) greater than 3, we calculate the SFR from (Kennicutt, 1998), as derived from the 3 GHz radio emission. This is done using the infrared radio correlation (IRRC, , Delvecchio et al., 2021) and a spectral index of = -0.8. The high spatial resolution (0.75″) of the VLA 3 GHz observations helps mitigate blending issues in crowded regions, which can affect FIR flux measurements. For sources with SNR3GHz ¡ 3, we compute from 24m fluxes using the conversion from Wuyts et al. (2008). The results are listed in Table. 1.
As shown in Fig. 1, SFRradio and SFR24μm are consistent. For galaxy 839791 in COSMOS-SBC6, which has a radio luminosity twice that of the crossover luminosity described in Wang et al. (2024), the threshold above which radio emission is dominated by AGN, we use SFR24μm to minimize the contamination from AGN-related radio emission.
In addition, we calculated SFRs by scaling the total derived from the best-fit SED of the integrated FIR photometry of each protocluster (Sillassen et al., 2024) to the dust continuum emission of individual members, assuming the same dust temperature for all members. We find that the scaled SFR is generally 1.6 times higher than SFRradio, this is more prominent for those galaxies with 100M⊙ yr-1, while galaxies with higher SFRradio have similar scaled SFR.
@@ 2.3 Gas masses
The molecular gas mass is converted from the Rayleigh-Jeans dust continuum at 850m, , as calibrated with the prescription in Hughes et al. (2017). This method leverages dust continuum emission, which is intrinsically associated with gas in the entire cold molecular gas reservoir (Genzel et al., 2015; Scoville et al., 2016), offering a much more sensitive measure than high- CO emission, which primarily traces the dense gas phase only (Liu et al., 2015; Daddi et al., 2015; Valentino et al., 2020a) and have limited detections due to the narrow line width compared to continuum observations. The dust continuum at 3 mm and 2 mm from the NOEMA and ALMA observations of the NICE survey, closely corresponding to rest-frame 850 m for the cluster galaxies, is used to derive gas masses. The NOEMA and ALMA observations have average angular resolutions of 4.1″and 0.7″, respectively. For spatially resolved sources, we use a Gaussian model to fit the visibilities. Unresolved sources are modeled as point sources.
We also estimated the gas masses from CO(3-2) or CO(4-3) line emission, , by converting them to the ground-state CO(1-0) luminosity, assuming the CO spectral line energy distribution (CO-SLED) of star-forming galaxies at = 2.0 - 2.7 from the ASPECS survey (Boogaard et al., 2020). These galaxies exhibit slightly higher CO excitation compared to main-sequence galaxies at ¡ 2 (Valentino et al., 2020b), but overall agree, within uncertainties, with the average SLEDs of SMGs at = 1.2 – 4.8 (Birkin et al., 2021). The conversion factor was derived using the mass–SFR–metallicity relation from Genzel et al. (2015) and the metallicity dependence of from Tacconi et al. (2018) for our sample.
In Fig. 2, we compare the gas masses estimated using the two methods. The two methods yield consistent results within 0.3 dex, with , approximately 1.5 times higher than . A similar comparison of CO(1–0) and 1.2 mm dust continuum measurements in galaxies within the Spiderweb protocluster also shows comparable results, albeit with large scatters (Zhang et al., 2024). Table. 1 lists values and gas masses calculated with different tracers. In the following, we adopt gas masses estimated from dust continuum for consistency.
@@
Notes. Columns: (1)–(6) protocluster ID, galaxy ID in COSMOS2020 (Weaver et al., 2022), distance to protocluster center, redshift, stellar mass and SFR as given in Sillassen et al. (2024). (7) CO-to-H2 conversion factor. (8) & (9) Molecular gas masses from dust continuum and from CO. (10) SFR from 3GHz observations, with 3 upper limits for non-detections. (11) SFR from 24m fluxes adopting conversion factors from Wuyts et al. (2008). For details, see Section 2. ∗: This source could be a blend of several galaxies (Sillassen et al., 2022). ∗∗: Radio-loud galaxy.
@@ 3.1 Enhanced Star Formation in Massive Halos in the Early Universe
in clusters increases with redshift, following an empirical relation of (1+)n (see Alberts & Noble 2022 for a review). The NICE survey extends such studies to ¿2. We adopt the integrated SFR and best estimates of halo masses of the eight protoclusters from (Sillassen et al., 2024) to obtain the total SFR per unit halo mass (). The SFR is from the FIR spectral energy distribution (SED) fitting using the dust templates in Magdis et al. (2012) and the halo masses are the average results from six methods including stellar mass to halo mass relations, overdensity with galaxy bias, and NFW profile fitting to radial stellar mass density. In Fig. 3, we included the eight massive protoclusters in NICE-COSMOS, and LH-SBC3 at = 3.95, the most distant protocluster identified in the NICE survey (Zhou et al., 2024), along with other protoclusters: RO1001 at = 2.91 (Kalita et al., 2021) and CL J1001 at = 2.51 (Wang et al., 2018), which are prototypes within the NICE survey, and SPT2349-56 at = 4.3 (Hill et al., 2022) and Distant Red Core at = 4.0 (Long et al., 2020), which are starbursting protoclusters similar to the NICE clusters. We found that increases by an order of magnitude from = 2 to = 4, following the evolutionary trend observed in galaxy clusters up to = 1 ( (1+)5.4, Webb et al., 2013), which has recently been shown to extend to clusters at = 1.6–2 (Smail, 2024). These values are one to two orders of magnitude higher than those observed in the field ( (1+)3-4 Sargent et al., 2012; Ilbert et al., 2013, 2015). This may be due to the evolution of the SFMS, modulated by the evolution of the relation, and the reduction of the quenched fraction (Elias et al., in prep.).
3.2 Star formation main sequence in dense environments

### v3

@@
Massive protoclusters at , the peak of the cosmic star formation history, are key to understanding the formation mechanisms of massive galaxies in today’s clusters. However, studies of protoclusters at these high redshifts remain limited, primarily due to small sample sizes and heterogeneous selection criteria. For this work, we conducted a systematic investigation of the star formation and cold gas properties of member galaxies of eight massive protoclusters in the COSMOS field , using the statistical and homogeneously selected sample from the Noema formIng Cluster survEy (NICE). Our analysis reveals a steep increase in the star formation rates per halo mass () with redshifts in these intensively star-forming protoclusters, reaching values one to two orders of magnitude higher than those observed in the field at ¿ 2. We further show that instead of an enhancement of starbursts, this increase is largely driven by the concentration of massive and gas-rich star-forming galaxies in the protocluster cores. The member galaxies still generally follow the same star-forming main sequence as in the field, with a moderate enhancement at the low-mass end. Notably, the most massive protocluster galaxies ( ¿ 81010M⊙) exhibit higher and than their field counterparts, while remaining on the star-forming main sequence. These gas-rich, massive, and star-forming galaxies are predominantly concentrated in the protocluster cores and are likely progenitors of massive ellipticals in the center of today’s clusters. These results suggest that the formation of massive galaxies in such environments is sustained by substantial gas reservoirs, which in turn support persistent star formation and drive early mass assembly in forming cluster cores.
@@
The environmental influence on the progenitors of massive galaxies in local clusters is critical for understanding the formation of these galaxies (Peng et al., 2010; Behroozi et al., 2019; van der Burg et al., 2020) . Over the past decade, extensive studies of massive protoclusters at the peak of the cosmic star formation history (), selected through various methods, have been inconclusive in producing an unambiguous consistent picture of the environmental effects on star formation and cold gas properties of galaxies in these structures (e.g., Rudnick et al., 2017; Gómez-Guijarro et al., 2019; Jin et al., 2023, 2024; Bakx et al., 2024; Pensabene et al., 2024). Specifically, some structures show a prevalence of submillimeter galaxies (SMGs) or ultra-luminous infrared galaxies (LIRGs), suggesting enhanced star formation in their galaxies as compared to average field environments (e.g., Casey, 2016; Wang et al., 2016; Oteo et al., 2018). Other studies, primarily based on H emitters in protoclusters, report little to no difference in the star-forming main sequence (SFMS) compared to their field counterparts (e.g, Valentino et al., 2015; Shimakawa et al., 2018; Pérez-Mart´ınez et al., 2023). Molecular gas is the fuel for star formation. The well-established star formation law (Kennicutt, 1998; Gao & Solomon, 2004) highlights the role of internal secular processes in the conversion of cold gas into stars. Additionally, a bimodal scenario has been proposed to describe the observed offset between starburst galaxies and normal star-forming galaxies, where starburst galaxies show higher star formation efficiency (SFE), which may result from a higher merger rate or a more compact distribution of molecular gas, among other factors (Daddi et al., 2010; Sargent et al., 2014; Liu et al., 2019).
@@
In dense environments, increased interactions can compress gas, thus enhancing SFEs, while cold gas inflows through filaments connected to protoclusters can replenish gas reservoirs, thus sustaining high gas fractions. Deep CO observations of cluster galaxies 1.6 reveal a consistent or enhanced gas fraction compared to the field counterparts (Rudnick et al., 2017; Noble et al., 2017). In the Spiderweb protocluster at =2.15, Dannerbauer et al. (2017) identified a massive and extended CO(1-0) disk around the central radio galaxies. Follow-up analysis by Pérez-Mart´ınez et al. (2025) revealed a decreasing gas fraction with increasing stellar mass, showing a particularly steep decline above log(/M⊙) = 10.5. In CL J1001, a starbusting protocluster at = 2.51, Wang et al. (2018) observed decreasing gas fractions toward the cluster center, suggesting the imminent formation of a passive core by 2. Tadaki et al. (2019) found a declining gas fraction and gas depletion time with increasing stellar mass in galaxies in three protoclusters traced by radio galaxies at , and suggested that gas accretion is accelerated in less massive galaxies and suppressed in more massive ones. At even higher redshifts ( 4), extreme starbusting protocluster cores such as SPT2349-56 and Distant Red Core (DRC) exhibit lower gas fractions compared to field SMGs (Long et al., 2020; Hill et al., 2022). Nevertheless, there is a paucity of protoclusters at high redshift with deep multi-wavelength coverage to comprehensively link the stellar and gas properties of member galaxies. Such observations are essential for revealing the physical mechanisms driving the formation of massive galaxies in these dense environments.
@@
The Noema formIng Cluster survEy (NICE, Zhou et al. 2024) is a 159-hour NOEMA Large Program (ID: M21AA, PIs: E. Daddi and T. Wang) targeting 48 massive protocluster candidates across multiple fields, complemented by a 40-hour ALMA program (ID: 2021.1.00815.S) focusing on 25 southern candidates. The primary goal of these programs is to spectroscopically confirm protoclusters at cosmic noon through CO emission from their member galaxies, significantly expanding the sample size in this field of research. Notably, eight protoclusters at 1.5 ¡ ¡ 4 have been confirmed in the COSMOS field (Sillassen et al., 2024). The COSMOS field provides extensive ancillary data across optical, near-infrared (NIR), far-infrared (FIR), submillimeter, and radio wavelengths (Weaver et al., 2022; Jin et al., 2018; Smolčić et al., 2017), which, combined with the CO observations from the NICE survey, offer a unique opportunity to investigate the environmental impact on galaxy evolution.
@@
For this paper we investigated the star formation activity and cold gas content of eight massive protoclusters in the COSMOS field, identified through the NICE survey (hereafter NICE-COSMOS). Our analysis focuses on the protoclusters’ central region (within one virial radius), as detailed in the following sections. We adopt a standard CDM cosmology with = 70 km s-1 Mpc-1, = 0.3, and = 0.7, and a Chabrier (2003) initial mass function.
@@ 2.1 NICE-COSMOS protoclusters
The eight protoclusters in NICE-COSMOS, along with their member identification, are detailed in Sillassen et al. (2024). Briefly, the members are selected with — - — ¡ 0.1(1 + ) within a virial projected radius based on the COSMOS2020 catalog (Weaver et al., 2022). The NICE survey selects intensively star-forming massive protoclusters at 2 4, which are chosen as overdensities of high- massive galaxies traced by red IRAC sources in association with SPIRE-350m peakers that show intense star formation. We list the selection criteria here:
@@
, , and are the Herschel/SPIRE fluxes at 250m, 350m, and 500m.
Figure 8 shows the spatial distribution of all members including CO and continuum detections in the eight protoclusters in NICE-COSMOS, which are primarily concentrated at the cluster centers. For the following analysis, we focus on galaxies within one virial radius, 1, as defined in (Sillassen et al., 2024), as this area is also well covered by the NOEMA and ALMA observations. The virial radii range from 141 pkpc to 429 pkpc, and the number of member galaxies per structure varies between 13 and 31.
@@ 2.2 Stellar masses and star formation rates
We used stellar masses and star formation rates (SFRs) derived with LePhare (Arnouts et al., 2002) in COSMOS2020 (Weaver et al., 2022). For galaxies detected in the FIR, specifically those in the super-deblended catalog (Jin et al., 2018), we also explored various SFR tracers to obtain accurate measurements.
We first cross-matched the members in the super-deblended catalog (Jin et al., 2018) with a separation smaller than 1. The matches have a median separation of 0.15. Notably, the 3GHz and 24m fluxes are measured directly at the galaxy positions, hence the risk of misidentification or flux misattribution is substantially reduced. For galaxies detected at 3 GHz (Smolčić et al., 2017) with a signal-to-noise ratio (S/N) greater than 3, we calculated the SFR from (Kennicutt, 1998), as derived from the 3 GHz radio emission. This was done using the infrared radio correlation (IRRC, , Delvecchio et al., 2021) and a spectral index of = -0.8. The high spatial resolution (0.75″) of the VLA 3 GHz observations helps mitigate blending issues in crowded regions, which can affect FIR flux measurements. For sources with S/N3GHz ¡ 3, we computed from 24m fluxes using the conversion from Wuyts et al. (2008). The results are listed in Table 1.
As we show in Fig. 1, SFRradio and SFR24μm are consistent. For galaxy 839791 in COSMOS-SBC6, which has a radio luminosity twice that of the crossover luminosity described in Wang et al. (2024), the threshold above which radio emission is dominated by AGN, we used SFR24μm to minimize the contamination from AGN-related radio emission.
In addition, we calculated SFRs by scaling the total derived from the best-fit SED of the integrated FIR photometry of each protocluster (Sillassen et al., 2024) to the dust continuum emission of individual members, assuming the same dust temperature for all members. We find that the scaled SFR is generally 1.6 times higher than SFRradio; this is more prominent for those galaxies with 100M⊙ yr-1, while galaxies with higher SFRradio have similarly scaled SFRs.
@@ 2.3 Gas masses
The molecular gas mass is converted from the Rayleigh-Jeans dust continuum at 850m, , as calibrated with the prescription in Hughes et al. (2017). This method leverages dust continuum emission, which is intrinsically associated with gas in the entire cold molecular gas reservoir (Genzel et al., 2015; Scoville et al., 2016), and offers a much more sensitive measure than high- CO emission, which primarily traces the dense gas phase only (Liu et al., 2015; Daddi et al., 2015; Valentino et al., 2020a) and has limited detections due to the narrow line width compared to continuum observations. The dust continuum at 3 mm and 2 mm from the NOEMA and ALMA observations of the NICE survey, closely corresponding to rest-frame 850 m for the cluster galaxies, was used to derive gas masses. The NOEMA and ALMA observations have average angular resolutions of 4.1″and 0.7″, respectively. For spatially resolved sources, we used a Gaussian model to fit the visibilities. Unresolved sources were modeled as point sources.
We also estimated the gas masses from CO(3-2) or CO(4-3) line emission, , by converting them to the ground-state CO(1-0) luminosity, assuming the CO spectral line energy distribution (CO-SLED) of star-forming galaxies at = 2.0 - 2.7 from the ASPECS survey (Boogaard et al., 2020). These galaxies exhibit slightly higher CO excitation compared to main sequence galaxies at ¡ 2 (Valentino et al., 2020b), but overall agree, within the uncertainties, with the average SLEDs of SMGs at = 1.2 – 4.8 (Birkin et al., 2021). The conversion factor was derived using the mass–SFR–metallicity relation from Genzel et al. (2015) and the metallicity dependence of from Tacconi et al. (2018) for our sample.
In Fig. 2 we compare the gas masses estimated using the two methods. The two methods yield consistent results within 0.3 dex, with approximately 1.5 times higher than . A similar comparison of CO(1–0) and 1.2 mm dust continuum measurements in galaxies within the Spiderweb protocluster also shows comparable results, albeit with large scatters (Zhang et al., 2024). Table 1 lists the values and gas masses calculated with different tracers. For consistency, in the following we adopt gas masses estimated from the dust continuum.
@@
Notes. Columns: (1)–(6) protocluster ID, galaxy ID in COSMOS2020 (Weaver et al., 2022), distance to protocluster center, redshift, stellar mass, and SFR as given in Sillassen et al. (2024). (7) CO-to-H2 conversion factor. (8) & (9) Molecular gas masses from dust continuum and from CO. (10) SFR from 3GHz observations, with 3 upper limits for nondetections. (11) SFR from 24m fluxes adopting conversion factors from Wuyts et al. (2008). For details, see Section 2. ∗: This source could be a blend of several galaxies (Sillassen et al., 2022). ∗∗: Radio-loud galaxy.
@@ 3.1 Enhanced Star Formation in Massive Halos in the Early Universe
The value of in clusters increases with redshift, following an empirical relation of (1+)n (see Alberts & Noble 2022 for a review). The NICE survey extends such studies to ¿2. We adopted the integrated SFR and best estimates of halo masses of the eight protoclusters from (Sillassen et al., 2024) to obtain the total SFR per unit halo mass (). The SFR is from the FIR spectral energy distribution (SED) fitting using the dust templates in Magdis et al. (2012), and the halo masses are the average results from six methods including stellar mass-to-halo mass relations, overdensity with galaxy bias, and NFW profile fitting to radial stellar mass density. In Fig. 3 we include the eight massive protoclusters in NICE-COSMOS; LH-SBC3 at = 3.95, which is the most distant protocluster identified in the NICE survey (Zhou et al., 2024); protoclusters RO1001 at = 2.91 (Kalita et al., 2021) and CL J1001 at = 2.51 (Wang et al., 2018), which are prototypes within the NICE survey; and SPT2349-56 at = 4.3 (Hill et al., 2022) and Distant Red Core at = 4.0 (Long et al., 2020), which are starbursting protoclusters similar to the NICE clusters. We found that increases by an order of magnitude from = 2 to = 4, following the evolutionary trend observed in galaxy clusters up to = 1 ( (1+)5.4, Webb et al., 2013), which has recently been shown to extend to clusters at = 1.6–2 (Smail2024). These values are one to two orders of magnitude higher than those observed in the field ( (1+)3-4 Sargent et al., 2012; Ilbert et al., 2013, 2015). This may be due to the evolution of the SFMS, modulated by the evolution of the relation, and the reduction of the quenched fraction (Elias et al. in prep.).
3.2 Star-forming main sequence in dense environments